### v2

# Quantum advantage in batteries for Sachdev-Ye-Kitaev interactions

Abstract
A quantum advantage can be achieved in the unitary charging of quantum batteries if their cells are interacting. Here, we try to clarify with some analytical calculations whether and how this quantum advantage is achieved for sparse Sachdev-Ye-Kitaev interactions. By performing a simple modelization, we find that for $q$-point rescaled sparse SYK interactions the quantum advantage goes as $\Gamma\sim N^{\frac{\alpha}{q}-\frac{1}{2}}$, where $\alpha$ is related to the connectivity and $N$ is the number of cells.
I Introduction
Isolated quantum systems can be used to temporarily store some useful work as quantum batteries, which can be later used in a consumption center (see, e.g., Ref. [1] for a review). The extraction of the work can be performed by a cyclical change of the Hamiltonian parameters, and in this case the maximum average work extractable is called ergotropy [2]. In particular, by considering $N$ copies of the system, i.e., a quantum battery with $N$ cells, the total ergotropy can be larger than $N$ times the ergotropy of the single cell [3]. Interestingly, this amount of ergotropy can be also obtained without generating entanglement in the work extraction time-evolution [4]. In particular, work extraction from finite quantum systems can be related to some genuine quantum features, such as correlations [5, 6, 7, 8] and coherence [9]. Furthermore, more general optimization of the work extraction protocol can be performed by taking in account also the work fluctuations [10, 11] with a quasiprobability distribution [12, 13, 14]. On the other hand, concerning the duration time of the protocol, the average power of a unitary charging of a battery with cells was originally investigated in Ref. [15], where the presence of interactions between the cells during the charging allows to achieve a so-called quantum advantage [16]. There are several proposals for the realization of quantum batteries, for instance by using many-body interactions in spin systems [17], with a cavity assisted charging [18, 19, 20, 21, 48, 23, 24], in disordered chains [25] and fermionic Sachdev-Ye-Kitaev (SYK) interactions [26, 27, 28, 29], to name a few. In particular, the SYK model [30, 31] has received a large interest in the last years (see, e.g., Ref. [32] for a review). Among its properties, it displays a resistivity that is linear with respect to the temperature, exhibiting a so-called Planckian transport [32, 33, 34], a duality to a two-dimensional nearly anti-de Sitter space [31, 35, 36, 37], and many others, attracting the attention of both condensed matter and high energy researchers. Concerning its experimental realization, there are already some proposals, e.g., in solid state physics (see, e.g., Ref. [38] for a review) and in cavity quantum electrodynamics platforms [39]. Several studies investigate the mesoscopic physics by SYK model (see, e.g., Refs. [40, 41, 42, 43, 44, 45, 46]). Nonequilibrium dynamics was also investigated through this model, e.g., the heating of black holes at short times before the thermalization with a colder bath [47, 48, 49], the current driven by a double contact setup [50], and also a quantum quench [51]. Other studies concern eternal traversable wormholes [52], the Bekenstein-Hawking entropy [53] and the existence of anomalous power laws in the temperature dependent conductance [54]. Furthermore, the sparse SYK model was also investigated (see, e.g., Refs. [55, 56, 57, 58, 59, 60]). Having discussed the importance of the SYK model, here, we aim to clarify whether and how a quantum advantage in the power charging of quantum batteries can be achieved by exploiting SYK interactions. In order to do this we introduce some preliminaries notions in Sec. II and the SYK model in Sec. III. Then, we proceed by introducing a simple model in Sec. IV, which shows same scaling properties of SYK interactions as discussed in Sec. V. Finally, we discuss further our results in Sec. VI
II Preliminaries
We focus on a quantum battery having $N$ cells (or copies) and Hamiltonian $H_{0}=\sum_{i=1}^{N}h_{i}$, where $h_{i}$ is the Hamiltonian of the single cell. The battery is initially prepared in the ground-state $|\psi(0)\rangle=|0\rangle^{\otimes N}$ having zero energy, i.e., $h_{i}|0\rangle=0$ for all $i=1,\ldots,N$. The charging process leads the system to the final state $|\psi(\tau)\rangle=U_{\tau,0}|\psi(0)\rangle$ where $U_{t,0}$ is the time-evolution unitary operator generated by the time-dependent Hamiltonian $H(t)$ with $t\in[0,\tau]$ such that $H(0)=H(\tau)=H_{0}$. The time-evolved state $|\psi(t)\rangle$ with $t\in[0,\tau]$ defines a curve $\mathcal{C}$ in the projective space having Fubini-Study length [61]
[…]
where $\Delta H^{2}(t)=\langle H^{2}(t)\rangle_{t}-\langle H(t)\rangle_{t}^{2}$ and the averages are calculated with respect to the state $|\psi(t)\rangle$, i.e., we defined $\langle X\rangle_{t}=\langle\psi(t)|X|\psi(t)\rangle$ for any operator $X$. In this paper, we focus on a charging process generated by performing two quantum quenches at the initial and final times $t=0,\tau$, e.g., as in Ref. [26]. Thus, the time-dependent Hamiltonian is defined such that $H(t)=H_{0}$ for $t=0,\tau$ and $H(t)=H_{1}$ for $t\in(0,\tau)$, with $[H_{1},H_{0}]\neq 0$ in order to achieve a non-trivial dynamics. In this case, the unitary time-evolution operator is $U_{t,0}=e^{-iH_{1}t}$ when $t\in(0,\tau)$ and the Fubini-Study length in Eq. (1) simplifies to
$l(\mathcal{C})=\Delta H_{1}\tau\,,$ (2)
where
$\Delta H_{1}^{2}=\langle H_{1}^{2}\rangle_{0}-\langle H_{1}\rangle_{0}^{2}\,.$ (3)
We note that $\Delta H_{1}^{2}$ is equal to the work variance of the first quench at $t=0$, i.e., it can be obtained as $\Delta H_{1}^{2}=-G^{\prime\prime}(0)$, where $G(u)=\ln\langle e^{iH_{1}u}\rangle_{0}$ is the cumulant generating function of the work done in the first quench. Thus, since $l(\mathcal{C})$ does not scale with the number $N$ of cells, from Eq. (2) the duration time $\tau$ scales with $N$ as $\tau\sim 1/\Delta H_{1}$ and its scaling is completely determined by the quench at $t=0$. To define a quantum advantage, we always choose $H_{1}$ such that the gap $\Delta E_{1}=E_{max}-E_{min}$ between its maximum and minimum eigenvalues, which are $E_{max}$ and $E_{min}$, is not larger than the gap $\Delta E_{0}$ between the maximum and the minimum eigenvalues of $H_{0}$. Then, by considering $H_{0}$ with a maximum eigenvalue that linearly scales with $N$, we get $\Delta E_{1}\leq\Delta E_{0}\sim N$. From the Bhatia–Davis inequality we get
$\Delta H_{1}^{2}\leq(E_{max}-\mu)(\mu-E_{min})\,,$ (4)
where $\mu=\langle H_{1}\rangle_{0}$, from which we get $\Delta H_{1}^{2}\lesssim N^{2}$ and the bound $\tau\gtrsim 1/N$. Let us examine two different situations. In the first one the duration time is denoted with $\tau^{\|}$ and the charging process is obtained with $H_{1}=\sum_{i=1}^{N}v_{i}$, where $v_{i}$ is a local operator of the i-th cell, so that there are no interactions among the cells. In this case, from Eq. (3) we deduce $\Delta H_{1}^{2}\sim N$, thus the charging time scales as $\tau^{\|}\sim 1/\sqrt{N}$. In contrast, in the second situation the duration time is denoted with $\tau^{\sharp}$ and interactions are allowed during the charging process, e.g., $H_{1}=\sum_{i_{1},i_{2},\ldots,i_{q}}v_{i_{1}}v_{i_{2}}\cdots v_{i_{q}}$, and we can get the scaling $\tau^{\sharp}\sim 1/N^{a}$, with $a\in(1/2,1]$. By recalling that the average power is $P=W/\tau$, where $W=\langle H_{0}\rangle_{\tau}$ is the average work, we get the quantum advantage
$\Gamma=\frac{P^{\sharp}}{P^{\|}}=\frac{\tau^{\|}}{\tau^{\sharp}}\,,$ (5)
which scales as $\Gamma\sim N^{a-1/2}$. Since $a\leq 1$, we get the bound
$\Gamma\lesssim N^{\frac{1}{2}}\,.$ (6)
III Model
We focus on a charging process performed through $q$-point sparse fermionic interactions, i.e., $H_{1}=H_{SYK,q}$ with
$H_{SYK,q}=i^{\frac{q}{2}}\sum_{1\leq i_{1}<i_{2}<\ldots<i_{q}\leq 2N}x_{i_{1}i_{2}\ldots i_{q}}j_{i_{1}i_{2}\ldots i_{q}}\gamma_{i_{1}}\gamma_{i_{2}}\cdots\gamma_{i_{q}}\,,$ (7)
where $\{\gamma_{i},\gamma_{j}\}=2\delta_{i,j}$, and the quenched disorder $j_{i_{1}i_{2}\ldots i_{q}}$ has a Gaussian distribution with zero mean and variance $j^{2}(q-1)!/N^{q-1}$, where $j\sim\mathcal{O}(1)$. We consider $x_{i_{1}i_{2}\ldots i_{q}}=1$ with probability $p$ and $x_{i_{1}i_{2}\ldots i_{q}}=0$ with probability $1-p$, defining the connectivity of the fermions. We define
$C_{q}=\sum_{1\leq i_{1}<i_{2}<\ldots<i_{q}\leq 2N}\overline{x_{i_{1}i_{2}\ldots i_{q}}}\sim pN^{q}\,,$ (8)
where the bar denotes the average over the disorder. For our purposes, we focus on
$C_{q}\sim N^{\alpha}\,,$ (9)
where $\alpha\leq q$. We note that the maximum eigenvalue of the Hamiltonian in Eq. (7) scales as $E_{max}\sim\sqrt{N}$, so that the bound of Eq. (6) cannot be saturated, since from Eq. (4) we get $\Delta H^{2}_{1}\lesssim N$, and thus $\Gamma\lesssim\mathcal{O}(1)$. Then, in order to get a quantum advantage we rescale the energy spectrum, i.e., we define
$H_{1}=\sqrt{N}H_{SYK,q}\,.$ (10)
For $\alpha=q$ we expect that the bound in Eq. (6) is saturated, since the interactions are fermionic and fully-connected. In the next sections, we aim to clarify the role of the interactions and connectivity through a suitable simplification of the Hamiltonian $H_{1}$.
IV Interactions
In order to perform some analytical calculations, we aim to modelize an interaction $H_{1}$ giving the average power $P^{\sharp}$ of the SYK interactions in Eq. (10). Thus, we consider
$H_{1}=(JV)^{k}\,,$ (11)
where $V$ is a certain operator and we introduced a coupling $J$ such that $\Delta E_{1}\sim N$ for any positive integer $k$. Since for $k=1$, the gap scales as $\Delta E_{1}\sim N$, we consider the coupling $J=N^{-\frac{k-1}{k}}$. In this case we get
$\langle e^{iuH_{1}}\rangle_{0}=\langle e^{iu(JV)^{k}}\rangle_{0}\,.$ (12)
Since we are interested only to the variance $\Delta H_{1}^{2}$, we aim to calculate Eq. (12) for $u\to 0$. We search a function $f(z)$ such that
[…]
as $u\to 0$, where $C$ is a stationary phase path so that the dominant contribution will be given by the stationary point of $f(z)/u+zx$. We get
$\frac{f(z)}{u}=c\left(\frac{z^{k}}{u}\right)^{\frac{1}{k-1}}\,,$ (14)
where $c$ is a number that does not depend on $u$. For instance, for $k=2$ Eq. (13) reduces to the Hubbard-Stratonovich transformation. Thus, we get
[…]
As $N\to\infty$, to do some calculation we consider the generating function
$\langle e^{izJV}\rangle_{0}\sim e^{iN^{\frac{1}{k}}g(z)}\,,$ (16)
where $g(z)$ is a certain function, such that $g^{\prime}(0)\sim\mathcal{O}(1)$, whereas higher derivatives can scale with $N$, e.g., $g^{\prime\prime}(0)\sim N^{\beta}$. We consider that the stationary point $z$ approaches to zero as $u\to 0$, thus we can consider the Taylor expansion $g(z)=g^{\prime}(0)z+g^{\prime\prime}(0)z^{2}/2+\cdots$ in the integral in Eq. (15). By calculating the stationary point up to $\mathcal{O}(u^{3})$ terms, we get
$\langle e^{iuH_{1}}\rangle_{0}\sim\exp\left(iu{g^{\prime}}^{k}(0)N+c^{\prime}u^{2}{g^{\prime}}^{k-1}(0)g^{\prime\prime}(0)N^{\frac{2k-1}{k}}+\mathcal{O}(u^{3})\right)\,,$ (17)
where $c^{\prime}$ is a number that does not depend on $u$ and different from $c$. Thus, from Eq. (17) we get
$\Delta H_{1}^{2}\sim N^{\frac{2k-1}{k}+\beta}$ (18)
from which, if $W\sim N$, the quantum advantage scales as
$\Gamma\sim N^{\frac{k-1}{2k}+\frac{\beta}{2}}\,.$ (19)
As we will discuss in the next section, the quantum advantage in Eq. (19) scales as the one achieved for $H_{1}$ in Eq. (10) with $q=2k$. Thus, $\beta$ is expected to be related to the connectivity, i.e., to $\alpha$ defined in Eq. (9).
V Connectivity
We consider an operator $V$ of the form
$V=\sum_{i,j}y_{ij}\lambda_{ij}v_{i}v_{j}+\sum_{i}\lambda_{i}w_{i}\,,$ (20)
where $w_{i}$ are local operators commutating with the local operators $v_{i}$, i.e., $[w_{i},v_{j}]=0$ for all $i$ and $j$, $\lambda_{i}$ and $\lambda_{ij}$ are couplings, and $y_{ij}=1$ with probability $p_{1}$ and $y_{ij}=0$ with probability $1-p_{1}$, where $p_{1}$ is such that $p_{1}^{k}\sim p$. If $\lambda_{ij}=0$, it is easy to see that $g^{\prime}(0)\sim\mathcal{O}(1)$ but $g^{\prime\prime}(0)\sim\mathcal{O}(J)$, thus $\beta=(k-1)/k$. Then, from Eq. (19) $\Gamma\sim\mathcal{O}(1)$ and there is no quantum advantage, as expected. Let us consider $\lambda_{ij}$ nonzero and random with a Gaussian distribution having zero mean and variance $\sigma^{2}$. We start to focus on $p_{1}=1$. If $v_{i}$ are bosonic operators, i.e., $[v_{i},v_{j}]=0$ for all $i$ and $j$, by averaging over the disorder we get
$\langle e^{izJV}\rangle_{0}\sim\langle e^{izJ\sum_{i}\lambda_{i}w_{i}}e^{-\frac{1}{2}z^{2}J^{2}\sigma^{2}\sum_{i,j}v_{i}^{2}v_{j}^{2}}\rangle_{0}$ (21)
If $\langle 0|w_{i}^{2}|0\rangle-\langle 0|w_{i}|0\rangle^{2}=0$ for all $i$, e.g., $w_{i}$ are c-numbers, from Eq. (21), we deduce that $g^{\prime\prime}(0)\sim\mathcal{O}(JN\sigma^{2})$. However, if $v_{i}$ are bosonic operators then $\lambda_{ij}\sim 1/N$ in order to have $\Delta E_{1}\sim N$, so that $\sigma^{2}\sim N^{-2}$ and $\beta=1/k-2$ and there is no quantum advantage at all, although $p_{1}=1$, since, from Eq. (19), $\Gamma\to 0$. In contrast, let us consider $p_{1}$ arbitrary and $v_{i}$ as fermionic operators. In this case $\lambda_{ij}\sim\mathcal{O}(1)$, i.e., $\sigma\sim\mathcal{O}(1)$, in order to have $\Delta E_{1}\sim N$. In order to connect this simplified model to $H_{1}$ in Eq. (10), we can focus on $H_{1}$ in Eq. (11) with
$V=i\sum_{i,j}y_{ij}\lambda_{ij}\gamma_{i}\gamma_{j}+wN\,,$ (22)
where $w$ is a c-number. There are two extreme cases: $y_{ij}=1$ for all $i$ and $j$, or $y_{ij}=1$ for $i=2l-1$ and $j=2l$ with $l=1,\ldots,N$ and $y_{ij}=0$ otherwise. In the first case $p_{1}=1$ and we get a full-connectivity. In this case, from Eq. (22) we already get $\Delta H_{1}^{2}\sim N^{2}$ for $k=2$. To explicitly prove it, we can focus on $\lambda_{ij}$ such that is nonzero only for odd $i$ and even $j$, so that from an explicit calculation of the average in Eq. (3) we get $\Delta H^{2}_{1}\geq\sigma^{4}(N-1)^{2}/4$ for $k=2$, thus we deduce $\beta=2-(2k-1)/k$. In the second case, which is equivalent to $p_{1}\sim 1/N$, we get $V=i\sum_{l}\lambda_{2l-1,2l}\gamma_{2l-1}\gamma_{2l}+wN$, and by averaging over the disorder we get
$\langle e^{izJV}\rangle_{0}\sim e^{izJwN}e^{-\frac{1}{2}z^{2}J^{2}\sigma^{2}N}\,,$ (23)
such that we get $g^{\prime\prime}(0)\sim\mathcal{O}(J\sigma^{2})$ and $\beta=1/k-1$. Then, from Eq. (19), $\Gamma\sim\mathcal{O}(1)$. To derive $\beta$ for an arbitrary $p_{1}$ we have to calculate the disorder average of $\langle e^{izJV}\rangle_{0}$ in general. We find $g^{\prime\prime}(0)\sim\mathcal{O}(p_{1}JN)$, thus by noting that $q=2k$, since $V$ is a quadratic form in the fermionic operators, and $p\sim N^{\alpha-q}$ from Eq. (9), we get
$\beta=\frac{2(\alpha-q+1)}{q}\,,$ (24)
where the exponent $\alpha$ characterizes the connectivity as Eq. (9). Thus, for $H_{1}$ in Eq. (10), by performing some algebraic manipulations, Eq. (19) simplifies to
$\Gamma\sim N^{\frac{\alpha}{q}-\frac{1}{2}}\,,$ (25)
which is our main result. In particular, we note that the bound in Eq. (6) is saturated for $\alpha=q$.
VI conclusion
Recently, the charging of quantum batteries has received a large attention from the scientific community. Here, we performed some analytical calculations with the aim to understand the charging of these quantum batteries through SYK interactions. Thus, we focused on a sparse SYK model, and we analyze the resulting quantum advantage. We determined the exact form of the quantum advantage scaling exponent, which is $z=\alpha/q-1/2$, where $\alpha$ is related to the connectivity, i.e., how much the interactions are sparse, and $q$ to the order of the interaction, i.e., the number of points involved. From this expression of $z$, we see that as $\alpha$ approaches to $q$, the scaling exponent reaches the maximum value of $1/2$, so that the quantum advantage scales as the square root of the number of cells. Interestingly, we can see how, while more connectivity enhances the charging performances, an increasing of $q$ reduces them. In conclusion, we hope that our results can be useful to clarify the charging of quantum batteries when SYK interactions are employed, at least from an theoretical point of view.