### v7

# Quantum advantage in batteries for Sachdev-Ye-Kitaev interactions

Abstract
A quantum advantage can be achieved in the unitary charging of quantum batteries if their cells are interacting. Here, we try to clarify with some analytical calculations whether and how this quantum advantage is achieved for sparse Sachdev-Ye-Kitaev (SYK) interactions and in general for fermionic interactions with disorder. To do this we perform a simple modelization of the interactions. In particular, we find that for $q$-point rescaled sparse SYK interactions the quantum advantage goes as $\Gamma\sim N^{\frac{\alpha-q}{2}+\frac{1}{2}}$ for $\alpha\geq q/2$ and $\Gamma\sim N^{\frac{1}{2}-\frac{\alpha}{2}}$ for $q/2>\alpha\geq 0$, where $\alpha$ is related to the connectivity and $N$ is the number of cells.
Introduction. Isolated quantum systems can be used to temporarily store some energy as quantum batteries, which can be later used to perform useful work in a consumption center (see, e.g., Ref. [1] for a review). The extraction of the work can be performed by a cyclical change of the Hamiltonian parameters, and in this case the maximum average work extractable is called ergotropy [2]. In particular, by considering $N$ copies of the system, i.e., a quantum battery with $N$ cells, the total ergotropy can be larger than $N$ times the ergotropy of the single cell [3]. Interestingly, this amount of ergotropy can be also obtained without generating entanglement in the work extraction time-evolution [4]. In particular, work extraction from finite quantum systems can be related to some genuine quantum features, such as correlations [5, 6, 7, 8] and coherence [9]. Furthermore, a more general optimization of the work extraction protocol can be performed by taking in account also the work fluctuations [10, 11] with a quasiprobability distribution [12, 13, 14]. The role of entanglement in charging quantum batteries was also investigated (see, e.g., Ref. [15]) and, recently, a quantum charging distance was introduced in Ref. [16]. Concerning the duration time of the protocol, the average power of a unitary charging of a battery with cells was originally investigated in Ref. [17], where the presence of interactions between the cells during the charging allows to achieve a so-called quantum advantage [18, 19]. There are several proposals for the realization of quantum batteries, for instance by using many-body interactions in spin systems [20], with a cavity assisted charging [21, 22, 23, 24, 50, 26, 27], in disordered chains [28] and fermionic Sachdev-Ye-Kitaev (SYK) interactions [29, 30, 31], to name a few. In particular, the SYK model [32, 33] has received a large interest in the last years (see, e.g., Ref. [34] for a review). Among its properties, it displays a resistivity that is linear with respect to the temperature, exhibiting a so-called Planckian transport [34, 35, 36], a duality to a two-dimensional nearly anti-de Sitter space [33, 37, 38, 39], and many others, attracting the attention of both condensed matter and high energy physicists. Concerning its experimental realization, there are already some proposals, e.g., in solid state physics (see, e.g., Ref. [40] for a review) and in cavity quantum electrodynamics platforms [41]. Several studies investigate the mesoscopic physics by SYK model (see, e.g., Refs. [42, 43, 44, 45, 46, 47, 48]). Nonequilibrium dynamics was also investigated through this model, e.g., the heating of black holes at short times before the thermalization with a colder bath [49, 50, 51], the current driven by a double contact setup [52], and also a quantum quench [53]. Other studies concern eternal traversable wormholes [54], the Bekenstein-Hawking entropy [55] and the existence of anomalous power laws in the temperature dependent conductance [56]. Furthermore, the sparse SYK model was also investigated (see, e.g., Refs. [57, 58, 59, 60, 61, 62]). Having discussed the importance of the SYK model, here, we aim to clarify whether and how a quantum advantage in the power charging of quantum batteries can be achieved by exploiting SYK interactions. Firstly, let us discuss the quantum advantage in general.
Quantum advantage. In order to derive the correct scaling of the quantum advantage, we focus on a quantum battery having $N$ cells (or copies) and Hamiltonian $H_{0}=\sum_{i=1}^{N}h_{i}$, where $h_{i}$ is the Hamiltonian of the single cell. The battery is initially prepared in the ground-state $|\psi(0)\rangle=|0\rangle^{\otimes N}$ having zero energy, i.e., $h_{i}|0\rangle=0$ for all $i=1,\ldots,N$. The charging process leads the system to the final state $|\psi(\tau)\rangle=U_{\tau,0}|\psi(0)\rangle$ where $U_{t,0}$ is the time-evolution unitary operator generated by the time-dependent Hamiltonian $H(t)$ with $t\in[0,\tau]$ such that $H(0)=H(\tau)=H_{0}$. The time-evolved state $|\psi(t)\rangle$ with $t\in[0,\tau]$ defines a curve $\mathcal{C}$ in the projective space having Fubini-Study length [63]
[…]
where $\Delta H^{2}(t)=\langle H^{2}(t)\rangle_{t}-\langle H(t)\rangle_{t}^{2}$ and the averages are calculated with respect to the state $|\psi(t)\rangle$, i.e., we defined $\langle X\rangle_{t}=\langle\psi(t)|X|\psi(t)\rangle$ for any operator $X$. In this letter, for simplicity we focus on a charging process generated by performing two quantum quenches at the initial and final times $t=0,\tau$. In detail, the time-dependent Hamiltonian is defined such that $H(t)=H_{0}$ for $t=0,\tau$ and $H(t)=H_{1}$ for $t\in(0,\tau)$, with $[H_{1},H_{0}]\neq 0$ in order to achieve a non-trivial dynamics. In this case, the unitary time-evolution operator is $U_{t,0}=e^{-iH_{1}t}$ when $t\in(0,\tau)$ and the Fubini-Study length in Eq. (1) simplifies to
$l(\mathcal{C})=\Delta H_{1}\tau\,,$ (2)
where
$\Delta H_{1}^{2}=\langle H_{1}^{2}\rangle_{0}-\langle H_{1}\rangle_{0}^{2}\,.$ (3)
We note that $\Delta H_{1}^{2}$ is equal to the work variance of the first quench at $t=0$, i.e., it can be obtained as $\Delta H_{1}^{2}=-G^{\prime\prime}(0)$, where $G(u)=\ln\langle e^{iH_{1}u}\rangle_{0}$ is the cumulant generating function of the work done in the first quench. It is well known that the Fubini-Study length corresponds to the length of the geodetic connecting two points in a hyper-dimensional sphere having unit radius, so that in our case $l(\mathcal{C})\leq\pi$ for any value of $N$. Then, $l(\mathcal{C})\sim\mathcal{O}(1)$ if the final and initial states do not become arbitrarily close in the limit $N\to\infty$ and $l(\mathcal{C})$ does not scale with the number $N$ of cells. Therefore, from Eq. (2) the duration time $\tau$ scales with $N$ as
$\tau\sim\frac{1}{\Delta H_{1}}$ (4)
and its scaling is completely determined by the quench at $t=0$. To define a quantum advantage, as in Ref. [17] we always choose $H_{1}$ such that the gap $\Delta E_{1}=E_{max}-E_{min}$ between its maximum and minimum eigenvalues, which are $E_{max}$ and $E_{min}$, is not larger than the gap $\Delta E_{0}$ between the maximum and the minimum eigenvalues of $H_{0}$. Then, by considering $H_{0}$ with a maximum eigenvalue that linearly scales with $N$, we get $\Delta E_{1}\leq\Delta E_{0}\sim N$. From the Bhatia–Davis inequality we get
$\Delta H_{1}^{2}\leq(E_{max}-\mu)(\mu-E_{min})\,,$ (5)
where $\mu=\langle H_{1}\rangle_{0}$, from which we deduce that when $\Delta H_{1}^{2}\sim N^{2a}$ we get $a\leq 1$, e.g., we can consider $E_{min}=0$, so that $E_{max}\sim N$ and $\mu$ scales at most linearly with $N$. Furthermore, from Eq. (4) the duration time scales as
$\tau\sim\frac{1}{N^{a}}\,.$ (6)
Let us examine two different situations. In the first one we perform a parallel charging where the duration time is denoted with $\tau^{\|}$ and the charging process is obtained with $H_{1}=\sum_{i=1}^{N}v_{i}$, where $v_{i}$ is a local operator of the i-th cell, so that there are no interactions among the cells. In this case, from Eq. (3) we deduce $\Delta H_{1}^{2}\sim N$, thus $a=1/2$ and the charging time scales as
$\tau^{\|}\sim\frac{1}{\sqrt{N}}\,,$ (7)
so that it goes to zero as $N\to\infty$. Then, the average power is $P^{\|}=W^{\|}/\tau^{\|}$, where $W^{\|}=\langle H_{0}\rangle_{\tau^{\|}}$ is the average work. When the work scales as $W^{\|}\sim N$, we get the correct scaling
$P^{\|}\sim N^{\frac{3}{2}}\,.$ (8)
We note that a wrong result $P^{\|}\sim N$ follows by assuming that the duration time scales as the duration time of the charging of a single cell. In contrast, in the second situation the duration time is denoted with $\tau^{\sharp}$ and interactions are allowed during the charging process, e.g., $H_{1}=\sum_{i_{1},i_{2},\ldots,i_{q}}v_{i_{1}}v_{i_{2}}\cdots v_{i_{q}}$, and depending on the interaction we can get the scaling $\tau^{\sharp}\sim 1/N^{a}$ for some $a\in(1/2,1]$ as we will see for the rescaled SYK case. In this case the average power is $P^{\sharp}=W^{\sharp}/\tau^{\sharp}$, and by considering a parallel charging with an average work $W^{\|}=W^{\sharp}$, we get the quantum advantage
$\Gamma=\frac{P^{\sharp}}{P^{\|}}=\frac{\tau^{\|}}{\tau^{\sharp}}\,,$ (9)
which scales as
$\Gamma\sim N^{a-\frac{1}{2}}.$ (10)
In particular, since $a\leq 1$, we get the general bound
$\Gamma\lesssim N^{\frac{1}{2}}\,,$ (11)
which is difficult to saturate. However, we will show how it can be saturated by exploiting a system of fermions with quenched disorder.
Model. Let us investigate a charging process with many-body interactions that gives $P^{\sharp}$. We start to focus on a charging process performed through $q$-point sparse fermionic interactions, i.e., $H_{1}=H_{SYK,q}$ with
$H_{SYK,q}=i^{\frac{q}{2}}\sum_{1\leq i_{1}<i_{2}<\ldots<i_{q}\leq 2N}x_{i_{1}i_{2}\ldots i_{q}}j_{i_{1}i_{2}\ldots i_{q}}\gamma_{i_{1}}\gamma_{i_{2}}\cdots\gamma_{i_{q}}\,,$ (12)
where $q$ is an even positive integer, $\{\gamma_{i},\gamma_{j}\}=2\delta_{i,j}$, and the quenched disorder $j_{i_{1}i_{2}\ldots i_{q}}$ has a Gaussian distribution with zero mean and variance $j^{2}(q-1)!/N^{q-1}$, where $j\sim\mathcal{O}(1)$. We consider $x_{i_{1}i_{2}\ldots i_{q}}=1$ with probability $p$ and $x_{i_{1}i_{2}\ldots i_{q}}=0$ with probability $1-p$, defining the connectivity of the fermions. We define the average number of $q$-point connections
$C_{q}=\sum_{1\leq i_{1}<i_{2}<\ldots<i_{q}\leq 2N}\overline{x_{i_{1}i_{2}\ldots i_{q}}}\sim pN^{q}\,,$ (13)
where the bar denotes the average over the disorder. For our purposes, we focus on
$C_{q}\sim N^{\alpha}\,,$ (14)
where $\alpha\leq q$. We note that the maximum eigenvalue of the Hamiltonian in Eq. (12) scales as $E_{max}\sim\sqrt{N}$ for $\alpha=q$, so that the bound of Eq. (11) cannot be saturated, since from Eq. (5) we get at most $\Delta H^{2}_{1}\sim N$, and thus $a=1/2$ and, from Eq. (10), $\Gamma\sim\mathcal{O}(1)$. In particular, numerical investigations done for $q=4$ in Ref. [29] indicate that the power scales as Eq. (8). It is easy to see that under a rescaling of the spectrum $H_{1}\mapsto MH_{1}$ we get $\Delta H_{1}\mapsto M\Delta H_{1}$, where $M$ is a positive constant. Then, in order to get a quantum advantage, i.e., $a>1/2$, we rescale the energy spectrum, i.e., we define
$H_{1}=MH_{SYK,q}\,.$ (15)
For $\alpha=q$, $M=\sqrt{N}$ in order to get $E_{max}\sim N$, and we expect that the bound in Eq. (11) is saturated, since the interactions are fermionic and fully-connected. In the following, we aim to clarify the role of the interactions and connectivity through a suitable simplification of the Hamiltonian $H_{1}$. We start to calculate how the gap $\Delta E_{1}$ scales with $N$ for $q=2$, so that from Eq. (15) we get $H_{1}=i\sum A_{ij}\gamma_{i}\gamma_{j}$. By randomly generating $x_{ij}$, the nonzero entries of the skew-symmetric matrix $A_{ij}$ essentially are uniformly distributed, thus the disorder average of $\Delta E_{1}$ is determined by only these matrices. Let us consider $C_{2}\sim N^{\alpha}$ as in Eq. (14). For $1\leq\alpha\leq 2$, all the $2N$ fermions appear in the Hamiltonian $H_{1}$, so that $\Delta E_{1}\sim N$. In contrast, for $0\leq\alpha\leq 1$, only $\mathcal{O}(N^{\alpha})$ fermions appear in the Hamiltonian $H_{1}$, and thus $\Delta E_{1}\sim N^{\alpha}$. Then, we introduce a simplified model for $q=2$ as $H_{1}=V$ with
$V=N^{x\theta(x)}\left(i\sum_{i,j}y_{ij}\lambda_{ij}\gamma_{i}\gamma_{j}+wN\right)$ (16)
and from $V$ we get $q$-point interactions by considering $V^{k}$ with $k=q/2$, so that the simplified Hamiltonian for arbitrary $q$ reads
$H_{1}=(JV)^{k}\,,$ (17)
where $\theta(x)$ is the Heaviside step function, $y_{ij}=1$ with probability $p_{1}$ and $y_{ij}=0$ with probability $1-p_{1}$, where $p_{1}$ is such that $p_{1}^{k}\sim p$, so that $\sum\overline{y_{i_{1}i_{2}}\cdots y_{i_{q-1}i_{q}}}\sim C_{q}$ and $x\in[-1,1]$ is such that $p_{1}\sim 1/N^{x+1}$, i.e., $x=1-2\alpha/q$. Furthermore, we introduced a c-number $w$ that scales as $p_{1}$, i.e., $w\sim p_{1}$, and a coupling $J$ such that $\Delta E_{1}\sim N$ for any positive integer $k$. Since for $k=1$, the gap scales as $\Delta E_{1}\sim N$, we consider the coupling $J=N^{-\frac{k-1}{k}}$. Let us focus on an arbitrary operator $V$, not necessarily of the form in Eq. (16). In this case we get
$\langle e^{iuH_{1}}\rangle_{0}=\langle e^{iu(JV)^{k}}\rangle_{0}\,.$ (18)
Since we are interested only to the variance $\Delta H_{1}^{2}$, we aim to calculate Eq. (18) for $u\to 0$. We search a function $f(z)$ such that
[…]
as $u\to 0$, where $C$ is a stationary phase path so that the dominant contribution will be given by the stationary point of $f(z)/u+zx$. We get
$\frac{f(z)}{u}=c\left(\frac{z^{k}}{u}\right)^{\frac{1}{k-1}}\,,$ (20)
where $c$ is a number that does not depend on $u$. For instance, for $k=2$ Eq. (19) reduces to the Hubbard-Stratonovich transformation. Thus, we get
[…]
As $N\to\infty$, to do some calculation we consider the generating function
$\langle e^{izJV}\rangle_{0}\sim e^{iN^{\frac{1}{k}}g(z)}\,,$ (22)
where $g(z)$ is a certain function. We consider that the stationary point $z$ approaches to zero as $u\to 0$, thus we can consider the Taylor expansion $g(z)=g^{\prime}(0)z+g^{\prime\prime}(0)z^{2}/2+\cdots$ in the integral in Eq. (21). By calculating the stationary point up to $\mathcal{O}(u^{3})$ terms, we get
$\langle e^{iuH_{1}}\rangle_{0}\sim\exp\left(iu{g^{\prime}}^{k}(0)N+c^{\prime}u^{2}{g^{\prime}}^{2k-2}(0)g^{\prime\prime}(0)N^{\frac{2k-1}{k}}+\mathcal{O}(u^{3})\right)\,,$ (23)
where $c^{\prime}$ is a number that does not depend on $u$ and different from $c$. Thus, from Eq. (23) we get
$\Delta H_{1}^{2}\sim{g^{\prime}}^{2k-2}(0)g^{\prime\prime}(0)N^{\frac{2k-1}{k}}\,,$ (24)
where for $V$ of Eq. (16) we get
$g^{\prime}(0)\sim N^{x\theta(x)}\left(\frac{1}{N}\sum_{i\neq j}y_{ij}\lambda_{ij}C_{ij}+w\right)$ (25)
and
$ig^{\prime\prime}(0)\sim N^{2x\theta(x)+1}\frac{J}{N^{2}}\sum_{i\neq j}\sum_{k\neq l}y_{ij}y_{kl}\lambda_{ij}\lambda_{kl}C_{il}C_{jk}$ (26)
with $C_{ij}=i\langle\gamma_{i}\gamma_{j}\rangle_{0}$. We get
$i{g^{\prime}}^{2k-2}(0)g^{\prime\prime}(0)$ $\sim$ $JN^{2kx\theta(x)+1}\sum_{n=0}^{2k-2}\binom{2k-2}{n}\Lambda_{n}w^{2k-2-n}$ (27)
$=$ $\sum_{n=0}^{2k-2}K_{n}\,,$
where
$\Lambda_{n}$ $=$ $\frac{1}{N^{n+2}}\sum_{i_{1}\neq j_{1}}\cdots\sum_{i_{n+2}\neq j_{n+2}}y_{i_{1}j_{1}}\cdots y_{i_{n+2}j_{n+2}}\lambda_{i_{1}j_{1}}\cdots\lambda_{i_{n+2}j_{n+2}}$ (28)
×$\times C_{i_{1}j_{1}}\cdots C_{i_{n}j_{n}}C_{i_{n+1}j_{n+2}}C_{i_{n+2}j_{n+1}}\,.$
By averaging over the disorder, if the moment $\overline{\lambda^{n+2}_{ij}}$ is nonzero we get $\overline{\Lambda}_{n}\sim p_{1}$ and the term in the sum scales as
$K_{n}\sim N^{qx\theta(x)+\frac{2}{q}+\frac{2(\alpha-q)(q-1-n)}{q}}$ (29)
with $n=0,\ldots,2k-2$. In general, we can get $\overline{\Lambda}_{n}\sim p_{1}^{m_{n}}$ with $m_{n}=1,\ldots,n+2$, so that
$K_{n}\sim N^{qx\theta(x)+\frac{2}{q}+\frac{2(\alpha-q)(q-2-n+m_{n})}{q}}\,.$ (30)
For instance, for $n=6$, if only $\overline{\lambda^{2}_{ij}}\neq 0$, from Eq. (28) we get $\overline{\Lambda}_{6}\sim\overline{\Lambda}_{0}^{4}\sim p_{1}^{4}$, whereas if only $\overline{\lambda^{2}_{ij}}\neq 0$ and $\overline{\lambda^{6}_{ij}}\neq 0$, we get $\overline{\Lambda}_{6}\sim\overline{\Lambda}_{4}\overline{\Lambda}_{0}\sim p_{1}^{2}$, then we can get only $m_{6}\leq 4$ and in general we deduce that only for $m_{n}\leq n/2+3/2$ we can get Eq. (30). However, if only $\overline{\lambda_{ij}}\neq 0$, we get $\overline{\Lambda}_{n}\sim p_{1}^{n+2}$ and thus $m_{n}=n+2$. Thus, we get the general expansion
$\Delta H_{1}^{2}\sim N^{qx\theta(x)+2}\sum_{n}\sum_{m_{n}}c_{nm_{n}}N^{\frac{2(\alpha-q)(q-2-n+m_{n})}{q}}\,.$ (31)
For the SYK interactions we can calculate $\Delta H_{1}^{2}$ directly from the average over the disorder of Eq. (3) by using the Wick’s theorem, e.g., from Eq. (12), the disorder average of $H_{SYK,q}^{2}$ reads
$H^{2}_{SYK,q}=(-1)^{\frac{q}{2}}\frac{pj^{2}(q-1)!}{N^{q-1}}\sum_{1\leq i_{1}<\ldots<i_{q}\leq 2N}\gamma_{i_{1}}\cdots\gamma_{i_{q}}\gamma_{i_{1}}\cdots\gamma_{i_{q}}$ (32)
and the disorder average of $H_{SYK,q}$ is zero, from which from Eq. (15) with $M=N^{\frac{qx\theta(x)}{2}+\frac{1}{2}}$ we easily obtain the term with $n-m_{n}=k-2$ in the sum of Eq. (31), which reads
$\Delta H_{1}^{2}\sim N^{qx\theta(x)+2+\alpha-q}\,.$ (33)
The probability distribution of $\lambda_{ij}$ can be such that only $\overline{\lambda_{ij}^{2}}$ is nonzero, so that $\overline{\Lambda}_{2l-2}\sim\overline{\Lambda}_{0}^{l}\sim p_{1}^{l}$ and thus $m_{2l-2}=l$ and $n-m_{n}=k-2$ is selected in the sum of Eq. (31). In contrast, for the disorder free case with a homogeneous $\lambda_{ij}$, from Eq. (28) we get $\overline{\Lambda}_{n}\sim p_{1}$ for any $n$, then we obtain the higher term with $m_{n}=1$ and $n=2k-2$ in the sum of Eq. (31), which reads
$\Delta H_{1}^{2}\sim N^{qx\theta(x)+2+2\frac{\alpha-q}{q}}\,.$ (34)
Thus, the quantum advantage can scale in different ways, which are
$\Gamma\sim N^{\frac{qx\theta(x)}{2}+\frac{1}{2}+\frac{(\alpha-q)r}{q}}$ (35)
with $r=1,\ldots,q$, where $r=q+m_{n}-n-2$, $n=0,\ldots,q-2$ and $m_{n}\leq n/2+3/2$ or $m_{n}=n+2$. Then, we get different classes of Hamiltonian operators with at most $q$-point interactions with disorder exhibiting the scalings of Eq. (35) and the rescaled SYK belongs to one of these classes. In particular, the Hamiltonian operators of Eq. (17) give the representative elements of the classes. We note that the classes are a consequence of a connectivity given by the product $x_{i_{1}\ldots i_{q}}=y_{i_{1}i_{2}}\cdots y_{i_{q-1}i_{q}}$ and interactions having an order lower than $q$ coming from the exponentiation of $V$. For the SYK class, for $x\in[-1,0]$ we get
$\Gamma\sim N^{\frac{\alpha-q}{2}+\frac{1}{2}}$ (36)
for $\alpha\geq q/2$. In particular, we note that, while for $p\sim 1/N$ we get $q-\alpha=1$ and there is no quantum advantage, the bound in Eq. (11) is saturated for $\alpha=q$. We note that, while $\lambda_{ij}\sim\mathcal{O}(1)$ for fermions in order to have $\Delta E_{1}\sim N$, if instead of $\gamma_{i}$ we have bosonic operators $v_{i}$, i.e., such that $[v_{i},v_{j}]=0$ for all $i$ and $j$, for full-connectivity $\lambda_{ij}\sim 1/N$ in order to have $\Delta E_{1}\sim N$. Thus, it is easy to see that $\Gamma\sim N^{-1/2}$ and there is no quantum advantage. On the other hand, for $x\in(0,1]$ we get
$\Gamma\sim N^{\frac{1}{2}-\frac{\alpha}{2}}$ (37)
for $q/2>\alpha\geq 0$. Then, the bound in Eq. (11) is saturated for $\alpha=0$, e.g., with only one $q$-point connection. In this case, in order to obtain a global unitary time-evolution we have to consider $q=2N$, e.g., as in the example investigated in Ref. [17].
Conclusion. Recently, the charging of quantum batteries has received a large attention from the scientific community. Here, we performed some analytical calculations with the aim to understand the charging of these quantum batteries through fermionic interactions with disorder, showing that there exist different scalings for the quantum advantage defining different classes of interactions. Thus, we focused on a sparse SYK model, and we analyzed the resulting quantum advantage, explaining its origin by using a simplified model. We determined the exact form of the quantum advantage scaling exponent, which is $z=(\alpha-q)/2+1/2$, for $\alpha\geq q/2$, where $\alpha$ is related to the connectivity, i.e., how much the interactions are sparse, and $q$ to the order of the interaction, i.e., the number of points involved. From the found expression of $z$, we see that as $\alpha$ approaches to $q$, the scaling exponent reaches the maximum value of $1/2$, so that the quantum advantage scales as the square root of the number of cells and we get the best possible performance for this charging process. Interestingly, we can see how, while more connectivity enhances the charging performance, an increasing of $q$ reduces it if $\alpha$ remains constant. In contrast, we get $z=1/2-\alpha/2$ for $q/2>\alpha\geq 0$, thus the scaling exponent reaches the maximum value of $1/2$ when the connectivity exponent decreases to zero, $\alpha=0$. However, in this case the unitary time-evolution tends to be global only if there are interactions with $q\sim 2N$. In conclusion, we hope that our results can be useful to clarify the charging of quantum batteries when SYK interactions are employed, at least from a theoretical point of view, and how improved performance can be achieved with many-body interactions. In particular, batteries made with fermions showing quenched disorder in the interactions seem to be indispensable in order to get the scaling achieved for the quantum advantage.